# Fair Play for Individuals, Foul Play for Groups? Auditing Anonymization’s Impact on ML Fairness

### v1

@@ 4.3 Detailed Analysis #2: ML Fairness Across Target Distribution Variations
Figure 3: Effect of target distribution variation (i.e., thresholded at deciles ranging from 10% to 90%) on anonymity techniques ($10$-anonymity, $2$-diversity, $0.5$-closeness) regarding group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. Results are presented for the Adult dataset, with gender serving as the protected attribute for fairness evaluation.
The results across all datasets and protected attributes (Figures 3,  1620) reveal consistent trends regarding the impact of anonymization techniques on fairness and utility metrics. Notably, the anonymized models tend to follow the same general shape as the baseline across varying deciles, indicating that the overall distributional patterns of fairness and utility metrics are preserved under anonymization. However, the magnitude of these metrics differs: anonymization consistently worsens group fairness metrics such as MAD, SPD, and EOD, reflecting an amplification of disparities between demographic groups. Conversely, individual fairness metrics, particularly ASF and NCF, generally benefit from anonymization, showing improved consistency in predictions for similar inputs. On the other hand, utility metrics such as ACC, F1-score, and ROC AUC are negatively impacted, with performance often falling below the baseline. These findings align with the conclusions drawn for RQ1 in Section 4.1, further underscoring the trade-offs between privacy, fairness, and utility introduced by anonymization techniques.
@@ 4.4 Detailed Analysis #3: Impact of Data Size on Fairness in ML
To address whether data fraction impacts fairness as formulated in RQ3 in Section 3.2, we analyze the behavior of anonymization techniques across subsampled data fractions. Specifically, we vary the data fraction from 10% to 100% of the original dataset, subsampling the data randomly at each fraction level. Figure 4 illustrates the results for the Adult dataset with gender as the protected attribute; additional results for other datasets and protected attributes are provided in Figures 2125 in Appendix B.
Figure 4: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, $2$-diversity, $0.5$-closeness) in terms of group fairness metrics (MAD, EOD, SPD), individual fairness metrics (ALF, ASF, NCF), and utility metrics (Accuracy, F1-score, ROC AUC) in ML. This analysis is performed using the Adult dataset, considering gender as the protected attribute for fairness evaluation.
Figures 42125 reveal that performance and fairness metrics under anonymization techniques follow a relatively stable pattern across different data fraction levels sampled from 10% to 100%. For instance, anonymization introduces consistent increases in group fairness disparities compared to the baseline, regardless of the dataset size. Similarly, we observe a positive effect of anonymization on individual fairness that remains stable across data sizes. Performance metrics improve gradually with increasing data, as expected. In conclusion, these findings suggest that the trade-offs between privacy, fairness, and utility are primarily driven by the anonymization techniques and their parameter settings, rather than by the scale of the dataset. The random sampling of fractions introduces minor variability but does not fundamentally alter the patterns observed.
@@ 4.5 Detailed Analysis #4: Comparison of ML Classifiers in Fairness Under Anonymization
From Figure 52630, results show that the trends observed with XGBoost in previous sections (RQ1RQ4) are consistent across other classifiers. Specifically, anonymization continues to negatively affect group fairness and to improve individual fairness. Moreover, as expected, utility degrades under anonymization, regardless of the classifier. However, the magnitude of degradation varies slightly. XGBoost and LGBM tend to retain higher performance than Random Forest and MLP, especially in terms of F1-score and ROC AUC. Although minor differences exist, such as XGBoost’s marginal advantage in utility and fairness stability, the median values for all fairness and utility metrics are nearly identical across classifiers. This visual and statistical consistency reinforces that our findings are not specific to XGBoost. Hence, the conclusions drawn from XGBoost experiments can be considered broadly representative, supporting the generalizability of our results across different learning algorithms.
@@ 4.6 General Findings and Practical Guidelines
Anonymity negatively impacts group fairness in ML: Anonymization methods generally degraded group fairness metrics (i.e., MAD, EOD, and SPD) as privacy parameters tightened. By coarsening quasi-identifiers and suppressing records, these anonymity-based methods distort the true joint distribution of $(A,X,Y)$, skew subgroup prevalences and error rates, and thereby amplify accuracy and positive-rate gaps between privileged and unprivileged groups.
@@
XGBoost findings generalize across ML classifiers: Fairness and utility trends observed with XGBoost are consistent across other classifiers, such as Random Forest and Neural Networks (MLP). While XGBoost often achieves slightly better utility, the broader patterns—negative group fairness impacts, positive individual fairness outcomes, and utility trade-offs—remain similar. This consistency supports the generalizability of conclusions drawn from XGBoost experiments.
@@
  • Avoid median splits when binarizing continuous target variable, as they tend to maximize group disparities. Evaluate multiple cutpoints across deciles. We found that thresholds below the $30$th or above the $70$th percentile yields lower group disparities; whereas median splits (4040404060606060%) tend to exacerbate them.
  • Interpret individual fairness improvements cautiously. Improvements in individual-fairness scores (ASF, NCF) are primarily due to feature homogenization and may not reflect “genuine improvements in equitable treatment”. Practitioners should therefore exercise caution and perform targeted case audits to avoid the risk of “fair washing/hacking” [1, 2, 31], for example.
@@ 5 Conclusion and Perspectives
This study systematically audits the tradeoff between anonymization techniques and fairness in ML. Through a comprehensive analysis, we evaluated the effects of three well-known anonymization methods ($k$-anonymity, $\ell$-diversity, and $t$-closeness) on group and individual fairness metrics, as well as utility metrics, across multiple datasets and ML classifiers. By addressing five key research questions, our findings highlight the inherent trade-offs between privacy, fairness, and utility in anonymized ML models. Overall, our results show that anonymization tends to negatively affect group fairness metrics, often exacerbating disparities between demographic groups as privacy constraints increase. In contrast, similarity-based individual fairness metrics tend to improve under stronger anonymization, driven by increased data homogeneity. While this effect aligns with phenomena observed in differential privacy [18], it stems from data structural smoothing rather than fairness-aware optimization, and must therefore be interpreted with caution. Importantly, we demonstrate that these findings are robust across different datasets, protected attributes, target distributions, data scales, suppression levels, and learning algorithms, confirming the generalizability of our conclusions. This work opens several promising research directions. A theoretical characterization of how anonymization influences fairness, e.g., through the lenses of causal modeling or information-theoretic frameworks, remains an open challenge. Future efforts could also focus on designing privacy mechanisms that jointly optimize for fairness constraints, or on adapting anonymization strategies to richer settings such as multi-class classification or regression.
@@ Appendix A Description of Datasets
  • Adult Dataset: The Adult dataset, obtained from the UCI ML repository [9] originates from the 1994 U.S. Census Bureau database. For this study, we use the Reconstructed Adult dataset as described in [16], where the target variable, income, is represented as a discrete value. This variable can be binarized into $>\mathsf{ThresholdIncome}$ or $\leq\mathsf{ThresholdIncome}$, where $\mathsf{ThresholdIncome}$ serves as a configurable threshold. By default, the threshold is set to the median income value; however, in Section 4.3, we vary this threshold to analyze the impact of target distribution shifts on fairness and utility. After preprocessing and data cleaning, the dataset consists of $n=45,849$ samples and $10$ discrete and categorical attributes. The protected attributes used for fairness evaluation are gender and race, while the target attribute is income.
  • Compas Dataset: The Compas dataset, curated by ProPublica [4], contains defendants from Broward County, Florida, screened between 2013 and 2014. We restrict to Black and White individuals who received a COMPAS risk score within 30 days of arrest. After preprocessing, the dataset comprises $n=5,278$ and five attributes: race, sex, age, priors_count, and days_b_screening_arrest. The target variable is the COMPAS risk score (v_decile_score), which consists of a rating of $1-10$; the higher the score, the more likely the defendant is to re-offend. This target variable can also be binarized into $>\mathsf{RiskScore}$ or $\leq\mathsf{RiskScore}$, where $\mathsf{RiskScore}$ serves as a configurable threshold. By default, the threshold is set to the median score value (i.e., 3); however, as in Section 4.3, we vary this threshold (1–10) to analyze the impact of target distribution shifts on fairness and utility. The protected attributes used for fairness evaluation are gender and race.
  • ACSIncome Dataset: The ACSIncome dataset, sourced from the U.S. Census Bureau’s American Community Survey (ACS), represents a geographically distributed sample of individuals across U.S. states. Similar to the Adult dataset, the target variable income is categorized as $>\mathsf{ThresholdIncome}$ or $\leq\mathsf{ThresholdIncome}$, with $\mathsf{ThresholdIncome}$ configurable to several representative income levels (median by default). For this study, we use the 2018 1-Year ACS Public Use Microdata Sample (survey_year="2018" and horizon="1-Year") across all U.S. states. The whole dataset contains $n=1,599,229$ data points with 10 discrete and categorical attributes. To reduce computational resource consumption, we randomly sample $10\%$ of the data. The protected attributes for fairness evaluation are gender and race (i.e., SEX, RAC1P), while the target attribute is income.
@@ B.2 Impact of Record Suppression Levels on Fairness in ML
Figure 11: Impact of suppression level in anonymization ($k$-anonymity, $\ell$-diversity, $t$-closeness) on group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. Results with the Adult dataset with gender as the protected attribute for fairness evaluation.
@@ B.4 Impact of Data Size on Fairness in ML
Figure 21: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, $2$-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the Adult dataset, considering race as the protected attribute for fairness evaluation.
Figure 22: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, $2$-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the Compas dataset, considering gender as the protected attribute for fairness evaluation.
Figure 23: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, $2$-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the Compas dataset, considering race as the protected attribute for fairness evaluation.
Figure 24: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, $2$-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the ACSIncome dataset, considering gender as the protected attribute for fairness evaluation.
Figure 25: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, $2$-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the ACSIncome dataset, considering race as the protected attribute for fairness evaluation.

### v2

@@ 4.3 Detailed Analysis #2: ML Fairness Across Target Distribution Variations
Figure 3: Effect of target distribution variation (i.e., thresholded at deciles ranging from 10% to 90%) on anonymity techniques ($10$-anonymity, 22-diversity, $0.5$-closeness) regarding group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. Results are presented for the Adult dataset, with gender serving as the protected attribute for fairness evaluation.
The results across all datasets and protected attributes (Figure 3 and Figures 1620) reveal consistent trends regarding the impact of anonymization techniques on fairness and utility metrics. Notably, the anonymized models tend to follow the same general shape as the baseline across varying deciles, indicating that the overall distributional patterns of fairness and utility metrics are preserved under anonymization. However, the magnitude of these metrics differs: anonymization consistently worsens group fairness metrics such as MAD, SPD, and EOD, reflecting an amplification of disparities between demographic groups. Conversely, individual fairness metrics, particularly ASF and NCF, generally benefit from anonymization, showing improved consistency in predictions for similar inputs. On the other hand, utility metrics such as ACC, F1-score, and ROC AUC are negatively impacted, with performance often falling below the baseline. These findings align with the conclusions drawn for RQ1 in Section 4.1, further underscoring the trade-offs between privacy, fairness, and utility introduced by anonymization techniques.
@@ 4.4 Detailed Analysis #3: Impact of Data Size on Fairness in ML
To address whether data fraction impacts fairness as formulated in RQ4 in Section 3.2, we analyze the behavior of anonymization techniques across subsampled data fractions. Specifically, we vary the data fraction from 10% to 100% of the original dataset, subsampling the data randomly at each fraction level. Figure 4 illustrates the results for the Adult dataset with gender as the protected attribute; additional results for other datasets and protected attributes are provided in Figures 21–25 in Appendix B.
Figure 4: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, 22-diversity, $0.5$-closeness) in terms of group fairness metrics (MAD, EOD, SPD), individual fairness metrics (ALF, ASF, NCF), and utility metrics (Accuracy, F1-score, ROC AUC) in ML. This analysis is performed using the Adult dataset, considering gender as the protected attribute for fairness evaluation.
Figure 4 and Figures 2125 reveal that performance and fairness metrics under anonymization techniques follow a relatively stable pattern across different data fraction levels sampled from 10% to 100%. For instance, anonymization introduces consistent increases in group fairness disparities compared to the baseline, regardless of the dataset size. Similarly, we observe a positive effect of anonymization on individual fairness that remains stable across data sizes. Performance metrics improve gradually with increasing data, as expected. In conclusion, these findings suggest that the trade-offs between privacy, fairness, and utility are primarily driven by the anonymization techniques and their parameter settings, rather than by the scale of the dataset. The random sampling of fractions introduces minor variability but does not fundamentally alter the patterns observed.
@@ 4.5 Detailed Analysis #4: Comparison of ML Classifiers in Fairness Under Anonymization
From Figure 5 and Figures 2630, results show that the trends observed with XGBoost in previous sections (RQ1RQ4) are consistent across other classifiers. Specifically, anonymization continues to negatively affect group fairness and to improve individual fairness. Moreover, as expected, utility degrades under anonymization, regardless of the classifier. However, the magnitude of degradation varies slightly. XGBoost and LGBM tend to retain higher performance than Random Forest and MLP, especially in terms of F1-score and ROC AUC. Although minor differences exist, such as XGBoost’s marginal advantage in utility and fairness stability, the median values for all fairness and utility metrics are nearly identical across classifiers. This visual and statistical consistency reinforces that our findings are not specific to XGBoost. Hence, the conclusions drawn from XGBoost experiments can be considered broadly representative, supporting the generalizability of our results across different learning algorithms.
@@ 4.6 General Findings and Practical Guidelines
Fairness vs. anonymization has non-monotonic behavior: While one might expect fairness to degrade monotonically with stricter privacy levels (i.e., higher $k$, $\ell$, or $t$), our findings consistently show non-linear behavior across datasets and metrics. This is due to the complex interplay between generalization, suppression, and model learning dynamics. Specifically, anonymization alters the statistical dependencies between features (i.e., both sensitive and non-sensitive), as well as between features and the target. Since generalization may be applied to any quasi-identifier, including the sensitive attribute, the relative influence of the sensitive feature on the model can increase or decrease in unpredictable ways compared to the non-anonymized case ($k=1$). These shifts can result in fluctuating fairness outcomes that do not follow a simple degradation trend.
Anonymity negatively impacts group fairness in ML: Anonymization methods generally degraded group fairness metrics (i.e., MAD, EOD, and SPD) as privacy parameters tightened. By coarsening quasi‑identifiers and suppressing records, these anonymity-based methods distort the true joint distribution of $(A,X,Y)$, skew subgroup prevalences and error rates, and thereby amplify accuracy and positive‑rate gaps between privileged and unprivileged groups.
@@
XGBoost findings generalize across ML classifiers: Fairness and utility trends observed with XGBoost are consistent across other classifiers, such as Random Forest and Neural Networks (MLP). While XGBoost often achieves slightly better utility, the broader patterns, i.e., negative group fairness impacts, positive individual fairness outcomes, and utility trade-offs, remain similar. This consistency supports the generalizability of conclusions drawn from XGBoost experiments.
@@
  • Avoid median splits when binarizing continuous target variable, as they tend to maximize group disparities. Evaluate multiple cutpoints across deciles. We found that thresholds below the $30$th or above the $70$th percentile yields lower group disparities; whereas median splits (40406060%) tend to exacerbate them.
  • Interpret individual fairness improvements cautiously. Improvements in individual‐fairness scores (ASF, NCF) are primarily due to feature homogenization and may not reflect “genuine improvements in equitable treatment”. Practitioners should therefore exercise caution and perform targeted case audits to avoid the risk of “fair washing/hacking” [1, 2, 33], for example.
@@ 5 Conclusion and Perspectives
This study systematically audits the tradeoff between anonymization techniques and fairness in ML. Through a comprehensive analysis, we evaluated the effects of three well-known anonymization methods ($k$-anonymity, $\ell$-diversity, and $t$-closeness) on group and individual fairness metrics, as well as utility metrics, across multiple datasets and ML classifiers. By addressing five key research questions, our findings highlight the inherent trade-offs between privacy, fairness, and utility in anonymized ML models. Overall, our results show that anonymization tends to negatively affect group fairness metrics, often exacerbating disparities between demographic groups as privacy constraints increase. In contrast, similarity-based individual fairness metrics tend to improve under stronger anonymization, driven by increased data homogeneity. While this effect aligns with phenomena observed in differential privacy [19], it stems from data structural smoothing rather than fairness-aware optimization, and must therefore be interpreted with caution. Importantly, we demonstrate that these findings are robust across different datasets, protected attributes, target distributions, data scales, suppression levels, and learning algorithms, confirming the generalizability of our conclusions. This work opens several promising research directions. A theoretical characterization of how anonymization influences fairness, e.g., through the lenses of causal modeling or information-theoretic frameworks, remains an open challenge. In particular, understanding the impact on individual fairness, where local input neighborhoods are disrupted by generalization and suppression, calls for a principled framework to describe how anonymization transforms the geometry of the feature space. Future efforts could also focus on designing privacy mechanisms that jointly optimize for fairness constraints, or on adapting anonymization strategies to richer settings such as multi-class classification or regression.
@@ Appendix A Description of Datasets
  • Adult Dataset: The Adult dataset, obtained from the UCI ML repository [10] originates from the 1994 U.S. Census Bureau database. For this study, we use the Reconstructed Adult dataset as described in [17], where the target variable, income, is represented as a discrete value. This variable can be binarized into $\mathsf{ThresholdIncome}$ or $\leq\mathsf{ThresholdIncome}$, where $\mathsf{ThresholdIncome}$ serves as a configurable threshold. By default, the threshold is set to the median income value; however, in Section 4.3, we vary this threshold to analyze the impact of target distribution shifts on fairness and utility. After preprocessing and data cleaning, the dataset consists of $n=45,849$ samples and $10$ discrete and categorical attributes. The protected attributes used for fairness evaluation are gender and race, while the target attribute is income.
  • Compas Dataset: The Compas dataset, curated by ProPublica [4], contains defendants from Broward County, Florida, screened between 2013 and 2014. We restrict to Black and White individuals who received a COMPAS risk score within 30 days of arrest. After preprocessing, the dataset comprises $n=5,278$ and five attributes: race, sex, age, priors_count, and days_b_screening_arrest. The target variable is the COMPAS risk score (v_decile_score), which consists of a rating of 1101-10; the higher the score, the more likely the defendant is to re-offend. This target variable can also be binarized into $\mathsf{RiskScore}$ or $\leq\mathsf{RiskScore}$, where $\mathsf{RiskScore}$ serves as a configurable threshold. By default, the threshold is set to the median score value (i.e., 3); however, as in Section 4.3, we vary this threshold (1–10) to analyze the impact of target distribution shifts on fairness and utility. The protected attributes used for fairness evaluation are gender and race.
  • ACSIncome Dataset: The ACSIncome dataset, sourced from the U.S. Census Bureau’s American Community Survey (ACS), represents a geographically distributed sample of individuals across U.S. states. Similar to the Adult dataset, the target variable income is categorized as $\mathsf{ThresholdIncome}$ or $\leq\mathsf{ThresholdIncome}$, with $\mathsf{ThresholdIncome}$ configurable to several representative income levels (median by default). For this study, we use the 2018 1-Year ACS Public Use Microdata Sample (survey_year="2018" and horizon="1-Year") across all U.S. states. The whole dataset contains $n=1,599,229$ data points with 10 discrete and categorical attributes. To reduce computational resource consumption, we randomly sample $10\%$ of the data. The protected attributes for fairness evaluation are gender and race (i.e., SEX, RAC1P), while the target attribute is income.
@@ B.2 Impact of Record Suppression Levels on Fairness in ML
Figure 11: Impact of suppression level in anonymization ($k$-anonymity, $\ell$-diversity, $t$-closeness) on group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. Results with the Adult dataset with race as the protected attribute for fairness evaluation.
@@ B.4 Impact of Data Size on Fairness in ML
Figure 21: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, 22-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the Adult dataset, considering race as the protected attribute for fairness evaluation.
Figure 22: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, 22-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the Compas dataset, considering gender as the protected attribute for fairness evaluation.
Figure 23: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, 22-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the Compas dataset, considering race as the protected attribute for fairness evaluation.
Figure 24: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, 22-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the ACSIncome dataset, considering gender as the protected attribute for fairness evaluation.
Figure 25: Effect of varying data fraction on the performance of anonymity techniques ($10$-anonymity, 22-diversity, $0.5$-closeness) in terms of group fairness (MAD, EOD, SPD), individual fairness (ALF, ASF, NCF), and utility (Accuracy, F1-score, ROC AUC) metrics in ML. This analysis is performed using the ACSIncome dataset, considering race as the protected attribute for fairness evaluation.